OpenStudy (anonymous):

A bowl contains 3 red, 8 blue, and 7 black beads. margaret randomly selects 3 beads. fine thhe probability of getting a red, ble, and black marble, in that order.

OpenStudy (anonymous):

\[\frac{3}{18}\times \frac{8}{17}\times \frac{7}{16}\] hope the method is clear